OpenStudy (anonymous):

use substitution to solve the system of equations. y=x+1 8x-4y=0

OpenStudy (anonymous):

y=x+1 --------(1) 8x-4y=0 --------(2) in the second equation (2) substitute the value of y from (1) 8x - 4(x+1) = 0 8x - 4x -4 = 0 4x = 4 x = 4/4 x = 1 now substitute x = 1 in equation (1) y = 1 + 1 y = 2 so x=1 and y=2 is the solution